OpenStudy (anonymous):

help pweeeeeeeeeease i have the answer but i need to know how to do this =/ In the period January 2000 to March 2000, stock for Sapient Corporation, a consulting company that helps businesses put their operations on the internet, dropped 51% and then rose by 49%. By how many percent did price change over the whole period?

OpenStudy (anonymous):

Let the price equal to $100. Now 51% percent of 100 is $51 $100 - $51 =$49 Now add 49 % of $49 - %24.01 to this amount $49 The new amount is $73.01 $100 (initial price) - $73.01 (final price) is $26.99 $26.99 is 26.99 % of %100. Answer is 26.99 drop

OpenStudy (anonymous):

damn thats kind of hard

OpenStudy (anonymous):

Starting at 100%, the stock dropped to 49% of its original value. It then increase by 49% of the 49%, i.e., 0.49*0.49 = 0.2401 or 24.01% of the original price. Thus in the end it was 49% + 24.01% or 73.01% of its original value. It was down 26.99%.

OpenStudy (anonymous):

Ti is the initial total value. Lets begin by applying the first decrease: \[T_i(1- 0.51)\]Then, apply the increase:\[T_i(1-0.51)(1+0.49)\]Then, we have that the total rate of change is that multiplication:\[(1-0.51)(1+0.49)=1-0.51+0.49-0.51 \times 0.49=1-0.2699\]Therefore, the percentage is 100*(0.2699) wich is approximately 27%, and it is a drop because of the minus sign.

OpenStudy (anonymous):

u made it kind of more complicated but thank u i got it now

OpenStudy (anonymous):

Oh really? Sorry, I thought was easier this way...

OpenStudy (anonymous):

hey grim why would u add the 24.01 to the 49% again?

OpenStudy (anonymous):

because that is the increase....